OpenStudy (anonymous):

Televisions and computer screens are usually advertised based on the length of their diagonals. If the height of a computer screen is 11 in. and the width is 14 in., what is the length of the diagonal? (Round to the nearest inch.)

OpenStudy (anonymous):

you use the Pythagorean theorem to solve this so 11^2 + 14^2 = c^2

OpenStudy (anonymous):

if you want the answer then its 121 + 196 = 317 anddd the answer is square root of 317... not sure what that is off the top of my head

OpenStudy (anonymous):

somewhere between 17 and 18

OpenStudy (anonymous):

ok thanks for the help...

OpenStudy (anonymous):

you are welcome :)